# Impact of Small-Scale Gravity Waves on Tracer Transport

### v1

@@ Abstract
Abstract
Small-scale gravity waves play a crucial role in atmospheric tracer transport, yet their effects remain unresolved in climate models and must be parameterized. This study investigates how gravity waves influence large-scale tracer distributions, utilizing a multiple-scale analysis to systematically identify the governing terms of gravity wave-induced tracer fluxes. The analysis reveals both leading-order and next-order impacts: the former being the inertia-gravity wave-induced tracer Stokes drift, which acts perpendicular to both the large-scale tracer gradient and the wave number, while the latter becomes significant at lower latitudes where Coriolis effects diminish. A numerical framework is developed to incorporate these fluxes into a gravity wave parameterization model, potentially enhancing climate model accuracy without requiring explicit resolution of small-scale wave dynamics. Model validation against high-resolution wave-resolving simulations confirms the effectiveness of this approach. By improving the representation of gravity wave-induced tracer transport, this research advances the accuracy of climate simulations, particularly in their depiction of microphysics and radiative processes.
@@ 1 Introduction
Accurately simulating the three-dimensional distribution of tracers is crucial for climate modeling, as these tracers have significant radiative effects on Earth’s climate. In the zonal mean, large-scale tracer transport—such as that of ozone and water vapor—is primarily governed by thermally and wave-driven global circulations, including the Brewer-Dobson circulation [6, 5, 8]. Small-scale processes, such as gravity waves and turbulence, indirectly influence tracer distributions by driving and modifying these circulations. However, they can also directly affect tracer transport and mixing. Since climate models do not resolve these small-scale processes, their effects must be represented through parameterization.
Previous studies have shown that gravity waves can significantly impact tracer transport [23, 19, 14]. Some parameterization models account for certain aspects of this influence, such as gravity-wave induced vertical mixing of tracers [10]. However, a systematic analysis of the governing equations is still lacking, leaving the full nature of gravity wave-induced tracer fluxes uncharacterized. A more refined parameterization is needed to capture these effects comprehensively.
@@
This paper is structured as follows. Section 2 derives the inertial and mid-frequency gravity wave-induced tracer transport using a multiple-scale analysis of the governing equations. Using a WKB ansatz, we obtain expressions for wave-induced tracer fluxes in terms of wave amplitudes. We then show how these fluxes can be computed from the phase-space wave action density, a quantity used in gravity wave ray tracing models. The most important equations are summarized in Section 3. Section 4.1 describes the extension of the parameterization model MS-GWaM [see e.g. 7, 20, 12] to include these fluxes and its coupling to an idealized flow solver. We validate the model by comparing its results to wave-resolving reference simulations. Finally, we summarize and conclude our results in Section 5.
@@ 2.1 Multiple-scale analysis of the governing equations
The advection of a unit-less passive tracer mixing ratio $\psi$ is given by
@@
The material derivative is defined as $\mathrm{D}_{t}=\partial_{t}+\mathbf{v}\cdot\nabla$, where $\mathbf{v}=(u,v,w)^{\mathrm{T}}$ with $u$, $v$, and $w$ representing the wind components in the $x$-, $y$-, and $z$-directions, respectively. The wind velocity $\mathbf{v}$, potential temperature $\theta$, and Exner pressure $\pi$ are governed by the equations of motion for a dry, inviscid atmosphere
@@
Here, $\mathbf{u}$ represents the horizontal wind components, $c_{p}$ and $c_{V}=c_{p}-R$ are the specific heat capacities at constant pressure and volume, respectively, $R$ is the ideal gas constant of dry air, $g$ is the gravitational acceleration, and $f$ is the constant Coriolis frequency.
@@
where $f_{0}=1$ serves as a non-dimensional placeholder for the Coriolis parameter.
@@
Characteristic tracer mixing ratio $\widetilde{\Psi}$ $1$
As solutions to Equations 6, 7, 8, 9 and 10, we consider a superposition of expansions in terms of the scale-separation parameter for the reference atmosphere, the large-scale flow, and $A$ waves described by the WKB ansatz. The variables thus take the form
The Exner pressure and potential temperature profiles of the reference atmosphere are denoted by $\overline{\Pi}^{(m)}$ and $\overline{\Theta}^{(m)}$, respectively, while large-scale variables are represented using angle brackets. The remaining variables, proportional to $e^{i\beta\phi_{\alpha}/\varepsilon}$, correspond to wave contributions, where $\beta=1$ represents a fundamental wave and $\beta=2,3,...$ denote its second, third, and higher harmonics. To emphasize the slow variation of the large-scale flow, wave amplitudes, and local wavenumbers $\beta\mathbf{k}_{\alpha}=\beta\nabla\phi_{\alpha}/\varepsilon=\beta\nabla_{\mathbf{X}}\phi_{\alpha}$ and frequencies $\beta\omega_{\alpha}=-\beta\partial_{t}\phi_{\alpha}/\varepsilon=-\beta\partial_{\mathrm{T}}\phi_{\alpha}$, we introduce the slow variables $(\mathbf{X},\mathrm{T})=\varepsilon(\mathbf{x},t)$. Additionally, the wave components of wind, potential temperature, and Exner pressure include a factor $e^{\gamma}$ to account for different wave amplitude cases. We distinguish three regimes: large-amplitude waves ($\gamma=0$), weakly nonlinear waves ($\gamma=1$), and quasi-linear waves ($\gamma=2$). Each regime requires a slightly different treatment. For $\gamma=0$, we assume monochromatic waves with $A=1$. In the weakly nonlinear case ($\gamma=1$), we allow for a spectrum of gravity waves ($A\geq 1$) but neglect wave-wave interactions affecting the gravity wave itself. Only the quasi-linear case does not require these previous assumptions. We impose no amplitude assumptions on the tracer field, allowing us to determine its magnitude from the governing equations.
@@
Before proceeding with the multiple-scale analysis of the tracer advection equation, we briefly outline how we handle the $\varepsilon$-expansion of the large-scale and background atmosphere variables and introduce a compact notation, which will be used throughout Section 2. Equations 11, 12, 13 and 14 present the full expansion of the large-scale and reference atmosphere variables in $\varepsilon$. In a rigorous asymptotic expansion of Equations 6, 7, 8, 9 and 10, we would account for the different orders of magnitude in $\varepsilon$ for these variables, as done in [1]. However, since the equations in this paper are designed for numerical implementation, we retain the full expansion rather than isolating specific orders in $\varepsilon$. This approach is necessary, because, in numerical models, the large-scale variables represent the full resolved flow without any expansion in $\varepsilon$. To streamline notation, we introduce a compact representation for the large-scale variables, exemplified here for the large-scale tracer
@@
$\sum_{m=0}^{\delta}\varepsilon^{m}\overline{\Pi}^{(m)}\equiv\overline{\underline{\Pi}}\quad\text{and}\quad\sum_{m=0}^{\delta}\varepsilon^{m}\overline{\Theta}^{(m)}\equiv\overline{\underline{\Theta}}\;,$ (16)
@@
We now proceed with the multiple-scale analysis of the tracer advection equation, aiming to identify the dominant terms that describe the impact of gravity waves on large-scale tracer transport. By substituting the decompositions from Equations 11 and 12 into Equation 6 and sorting terms by powers of $\varepsilon$ and the phase factor $e^{i\beta\phi_{\alpha}/\varepsilon}$, we find that the first nonzero tracer wave amplitude, $\hat{\Psi}_{\alpha\beta}^{(\gamma+1)}$, is related to the leading-order wind amplitude by
@@
$i\beta\hat{\omega}_{\alpha}\hat{\Psi}_{\alpha\beta}^{(\gamma+1)}=\hat{\mathbf{V}}_{\alpha\beta}^{(0)}\cdot\nabla_{\mathbf{X}}\left\langle{\underline{\Psi}}\right\rangle\;,$ (17)
@@
where $\hat{\omega}_{\alpha}=\omega_{\alpha}-\mathbf{k}_{h,\alpha}\cdot\left\langle{\underline{\mathbf{V}}}\right\rangle$ is the intrinsic frequency. The next-to-leading-order tracer amplitudes are given by the relation
These results also account for the factor that the leading-order large-scale vertical wind, $\left\langle{W}\right\rangle^{(0)}=0$, vanishes, as demonstrated in Achatz et al. [1]. To derive the result for $\gamma=1$, we neglect the triad terms in Equation 18 associated with wave-wave interactions. Using the large-scale terms and the relations above, we obtain the following equation for the large-scale tracer mixing ratio
$\begin{split}\varepsilon\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{V}}}\right\rangle\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Psi}}\right\rangle=-\frac{\varepsilon^{\gamma}}{2}\sum_{\alpha=1}^{A}\sum_{\beta=1}^{\infty}\sum_{m=0}^{\infty}\sum_{n=\gamma+1}^{\infty}&\varepsilon^{(m+n+1)}\mathfrak{R}\left(\hat{\mathbf{V}}_{\alpha\beta}^{(m)}\cdot\nabla_{\mathbf{X}}\hat{\Psi}_{\alpha\beta}^{(n)^{*}}\right)\\ &+\varepsilon^{(m+n)}\mathfrak{I}\left(\hat{\mathbf{V}}_{\alpha\beta}^{(m)}\cdot\beta\mathbf{k}_{\alpha}\hat{\Psi}_{\alpha\beta}^{(n)^{*}}\right)\;.\end{split}$ (19)
Additionally, using the wind amplitude relations derived from the Exner pressure equation (Equation 10)
$i\beta\mathbf{k}_{\alpha}\cdot\hat{\mathbf{V}}_{\alpha\beta}^{(0)}$ $=0\;,$ (20)
$i\beta\mathbf{k}_{\alpha}\cdot\hat{\mathbf{V}}_{\alpha\beta}^{(1)}$ $=-\frac{1}{\overline{P}}\nabla_{\mathbf{X}}\cdot\left(\overline{P}\hat{\mathbf{V}}_{\alpha\beta}^{(0)}\right)\;,$ (21)
$i\beta\mathbf{k}_{\alpha}\cdot\hat{\mathbf{V}}_{\alpha\beta}^{(2)}$ $=-\frac{1}{\overline{P}}\nabla_{\mathbf{X}}\cdot\left(\overline{P}\hat{\mathbf{V}}_{\alpha\beta}^{(1)}\right)\;,$ (22)
where $\overline{P}=\overline{\rho}\overline{\theta}$ is the mass-weighted potential temperature of the reference atmosphere, we find that Equation 22 is strictly valid for $\delta=0$, expect for a small correction that we ignore in the following analysis. Consequently, Equation 19 can be rewritten as
@@
$\begin{split}\varepsilon\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{V}}}\right\rangle\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Psi}}\right\rangle=\mathcal{Q}^{(0)}+\mathcal{Q}^{(1)}+O\left[\varepsilon^{(2\gamma+4)}\right]\;.\end{split}$ (23)
@@
The leading-order terms on the right-hand side of Equation 19, summarized in Equation 23 as $\mathcal{Q}^{(0)}$, are $\mathcal{O}\left(\varepsilon^{(2\gamma+1)}\right)$ smaller than the large-scale advection and can be rewritten as
@@
Here, we use the fact that higher harmonics with $\beta>1$ do not contribute, since for those, $\hat{\mathbf{V}}_{\alpha\beta}^{(0)}=0$ [1] and, by Equation 17, also $\hat{\Psi}_{\alpha\beta}^{(\gamma+1)}=0$. Similarly, all terms in Equation 19 that are $\mathcal{O}\left(\varepsilon^{(2\gamma+2)}\right)$ smaller than the large-scale advection are grouped into $\mathcal{Q}^{(1)}$ in Equation 23 and can be rewritten as
@@
With gravity wave parameterization in mind, it is convenient to express the terms in Equations 24 and 25 using variables commonly computed in parameterization models, such as the wave action density and wavenumbers. These formulation will be presented in the following sections.

### v2

@@ Abstract
Small-scale gravity waves, with horizontal wavelengths of up to a few hundred kilometers and vertical wavelengths of up to a few kilometers, play a crucial role in atmospheric tracer transport. However, their effects remain unresolved in climate models and must be parameterized. This study investigates how gravity waves influence large-scale tracer distributions, utilizing a multiple-scale analysis to systematically identify the governing terms of gravity wave-induced tracer fluxes. The analysis reveals both leading-order and next-order impacts: the former being the inertia-gravity wave-induced tracer Stokes drift, which acts perpendicular to both the large-scale tracer gradient and the wave number, while the latter becomes significant at lower latitudes where Coriolis effects diminish. A numerical framework is developed to incorporate these fluxes into a gravity wave parameterization model, potentially enhancing climate model accuracy without requiring explicit resolution of small-scale wave dynamics. Model validation against high-resolution wave-resolving simulations confirms the effectiveness of this approach. By improving the representation of gravity wave-induced tracer transport, this research advances the accuracy of climate simulations, particularly in their depiction of microphysics and radiative processes.
@@ 1 Introduction
Accurately simulating the three-dimensional distribution of tracers is crucial for climate modeling, as these tracers have significant radiative effects on Earth’s climate. In the zonal mean, large-scale tracer transport—such as that of ozone and water vapor—is primarily governed by thermally and wave-driven global circulations [7, 6, 9]. Small-scale processes, such as gravity waves and turbulence, indirectly influence tracer distributions by driving and modifying these circulations. However, they can also directly affect tracer transport and mixing. These gravity waves, with horizontal and vertical wavelengths of up to a few hundred and a few kilometers, respectively, remain unresolved in climate models and are small compared to synoptic-scale processes, which extend over several thousand kilometers. Since climate models do not resolve these small-scale processes, their effects must be represented through parameterization.
[…]
@@
This paper is structured as follows. Section˜2 derives the inertial and mid-frequency gravity wave-induced tracer transport using a multiple-scale analysis of the governing equations. Using a WKB ansatz, we obtain expressions for wave-induced tracer fluxes in terms of wave amplitudes. We then show how these fluxes can be computed from the phase-space wave action density, a quantity used in gravity wave ray tracing models. The most important equations are summarized in Section˜3. Section˜4 describes the extension of the parameterization model MS-GWaM [see e.g. 8, 22, 14] to include these fluxes and its coupling to an idealized flow solver. We validate the model by comparing its results to wave-resolving reference simulations. Finally, we summarize and conclude our results in Section˜5.
@@ 2.1 Multiple-scale analysis of the governing equations
The advection of a unit-less passive tracer mixing ratio $\psi$, neglecting diffusion, is given by
@@
The material derivative is defined as $\mathrm{D}_{t}=\partial_{t}+\mathbf{v}\cdot\nabla$, where $\mathbf{v}=(u,v,w)^{\mathrm{T}}$ with $u$, $v$, and $w$ representing the wind components in the $x$-, $y$-, and $z$-directions, respectively. The gradient operator is defined as $\nabla=\partial_{x}\mathbf{e}_{x}+\partial_{y}\mathbf{e}_{y}+\partial_{z}\mathbf{e}_{z}$ with unit vectors $\mathbf{e}_{x}$, $\mathbf{e}_{y}$, and $\mathbf{e}_{z}$. The wind velocity $\mathbf{v}$, potential temperature $\theta$, and Exner pressure $\pi$ are governed by the equations of motion for a dry, inviscid atmosphere [1]
@@
Here, $\mathbf{u}$ represents the horizontal wind components, $c_{p}$ and $c_{V}=c_{p}-R$ are the specific heat capacities at constant pressure and volume, respectively, $R$ is the ideal gas constant of dry air, $g$ is the gravitational acceleration, and $f$ is the constant Coriolis frequency. Furthermore, we define the horizontal gradient operator $\nabla_{h}=\partial_{x}\mathbf{e}_{x}+\partial_{y}\mathbf{e}_{y}$.
@@
$\varepsilon=\frac{\widetilde{U}}{f\widetilde{L}}=\frac{1}{f\widetilde{T}}=\mathcal{O}\left(10^{-1}\right)\ll 1\;,$ (6)
where $\widetilde{U}$, $\widetilde{L}$, $\widetilde{H}$, and $\widetilde{T}$ denote characteristic synoptic-scale horizontal velocity, length, height, and time scales, respectively (see Table˜1), while $\widetilde{L}_{w}$, $\widetilde{H}_{w}$, and $\widetilde{T}_{w}$ represent the characteristic length, height, and times scales of gravity waves. These synoptic scales provide the basis for non-dimensionalizing Equations˜1, 2, 3, 4 and 5, yielding the system
@@
where $f_{0}=1$ serves as a non-dimensional placeholder for the Coriolis parameter. Here, $\mathrm{D}_{t}=\partial_{t}+\mathbf{v}\cdot\nabla$ denotes the non-dimensional material derivative.
@@
Characteristic tracer mixing ratio $\widetilde{\Psi}$ 11
As solutions to Equations˜7, 8, 9, 10 and 11, we consider a superposition of expansions in terms of the scale-separation parameter for the reference atmosphere, the large-scale flow, and gravity waves described by the WKB ansatz. For clarity, we restrict the main presentation to a locally monochromatic, large-amplitude wave in a weakly stratified atmosphere ($\delta=1$). Here, large-amplitude means that the wave-induced wind field is of the same order of magnitude as the large-sacle wind field. A more general expansion, including a spectrum of gravity waves with higher harmonics under both strong and weak stratification, is provided in Appendix˜A. The variables thus take the form
$\psi(\mathbf{x},t)=$ $\sum_{m=0}^{\infty}\varepsilon^{m}\left\langle{\Psi}\right\rangle^{(m)}(\mathbf{X},\mathrm{T})+\mathfrak{R}\sum_{m=0}^{\infty}\varepsilon^{m}\hat{\Psi}^{(m)}(\mathbf{X},\mathrm{T})e^{i\phi(\mathbf{X},\mathrm{T})/\varepsilon}\;,$ (12)
$\mathbf{v}(\mathbf{x},t)=$ $\sum_{m=0}^{\infty}\varepsilon^{m}\left\langle{\mathbf{V}}\right\rangle^{(m)}(\mathbf{X},\mathrm{T})+\mathfrak{R}\sum_{m=0}^{\infty}\varepsilon^{m}\hat{\mathbf{V}}^{(m)}(\mathbf{X},\mathrm{T})e^{i\phi(\mathbf{X},\mathrm{T})/\varepsilon}\;,$ (13)
$\begin{split}\theta(\mathbf{x},t)=&\sum_{m=0}^{1}\varepsilon^{m}\overline{\Theta}^{(m)}(\mathrm{Z})+\varepsilon^{2}\sum_{m=0}^{\infty}\varepsilon^{m}\left\langle{\Theta}\right\rangle^{(m)}(\mathbf{X},\mathrm{T})+\varepsilon^{2}\mathfrak{R}\sum_{m=0}^{\infty}\varepsilon^{m}\hat{\Theta}^{(m)}(\mathbf{X},\mathrm{T})e^{i\phi(\mathbf{X},\mathrm{T})/\varepsilon}\;,\end{split}$ (14)
$\begin{split}\pi(\mathbf{x},t)=&\sum_{m=0}^{1}\varepsilon^{m}\overline{\Pi}^{(m)}(\mathrm{Z})+\varepsilon^{2}\sum_{m=0}^{\infty}\varepsilon^{m}\left\langle{\Pi}\right\rangle^{(m)}(\mathbf{X},\mathrm{T})+\varepsilon^{3}\mathfrak{R}\sum_{m=0}^{\infty}\varepsilon^{m}\hat{\Pi}^{(m)}(\mathbf{X},\mathrm{T})e^{i\phi(\mathbf{X},\mathrm{T})/\varepsilon}\;,\end{split}$ (15)
The Exner pressure and potential temperature profiles of the reference atmosphere are denoted by $\overline{\Pi}^{(m)}$ and $\overline{\Theta}^{(m)}$, respectively, while large-scale variables are represented using angle brackets. The remaining variables, proportional to $e^{i\phi/\varepsilon}$, correspond to wave contributions. To emphasize the slow variation of the large-scale flow, wave amplitudes, and local wavenumbers $\mathbf{k}=\nabla\phi/\varepsilon=\nabla_{\mathbf{X}}\phi$ and frequencies $\omega=-\partial_{t}\phi/\varepsilon=-\partial_{\mathrm{T}}\phi$, we introduce the slow variables $(\mathbf{X},\mathrm{T})=\varepsilon(\mathbf{x},t)$.
@@
Before proceeding with the multiple-scale analysis of the tracer advection equation, we briefly outline how we handle the $\varepsilon$-expansion of the large-scale and background atmosphere variables and introduce a compact notation, which will be used throughout Section˜2. Equations˜12, 13, 14 and 15 present the full expansion of the large-scale and reference atmosphere variables in $\varepsilon$. In a rigorous asymptotic expansion of Equations˜7, 8, 9, 10 and 11, we would account for the different orders of magnitude in $\varepsilon$ for these variables, as done in [2]. However, since the equations in this paper are designed for numerical implementation, we retain the full expansion rather than isolating specific orders in $\varepsilon$. This approach is necessary, because, in numerical models, the large-scale variables represent the full resolved flow without any expansion in $\varepsilon$. To streamline notation, we introduce a compact representation for the large-scale variables, exemplified here for the large-scale tracer
@@
$\sum_{m=0}^{1}\varepsilon^{m}\overline{\Pi}^{(m)}\equiv\overline{\underline{\Pi}}\quad\text{and}\quad\sum_{m=0}^{1}\varepsilon^{m}\overline{\Theta}^{(m)}\equiv\overline{\underline{\Theta}}\;,$ (17)
@@
We now proceed with the multiple-scale analysis of the tracer advection equation, aiming to identify the dominant terms that describe the impact of gravity waves on large-scale tracer transport. By substituting the decompositions from Equations˜12 and 13 into Equation˜7 and sorting terms by powers of $\varepsilon$ and the phase factor $e^{i\phi/\varepsilon}$, we find that the first nonzero tracer wave amplitude, $\hat{\Psi}^{(1)}$, is related to the leading-order wind wave amplitude by
@@
$i\hat{\omega}\hat{\Psi}^{(1)}=\hat{\mathbf{V}}^{(0)}\cdot\nabla_{\mathbf{X}}\left\langle{\underline{\Psi}}\right\rangle\;,$ (18)
@@
where $\hat{\omega}=\omega-\mathbf{k}_{h}\cdot\left\langle{\underline{\mathbf{V}}}\right\rangle$ is the intrinsic frequency. This result shows that the gradients in the large-scale tracer field induce oscillations in the leading-order tracer perturbation field that share the same frequency as the wind fluctuations. The next-to-leading-order tracer amplitudes are given by the relation
@@
$i\hat{\omega}\hat{\Psi}^{(2)}=\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)\hat{\Psi}^{(1)}+\hat{\mathbf{V}}^{(1)}\cdot\nabla_{\mathbf{X}}\left\langle{\underline{\Psi}}\right\rangle\;.$ (19)
These results also account for the fact that the leading-order large-scale vertical wind, $\left\langle{W}\right\rangle^{(0)}=0$, vanishes, as demonstrated in Achatz et al. [2]. Using the large-scale terms and the relations above, we obtain the following equation for the large-scale tracer mixing ratio
$\begin{split}\varepsilon\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{V}}}\right\rangle\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Psi}}\right\rangle=-\frac{1}{2}\sum_{m=0}^{\infty}\sum_{n=1}^{\infty}&\varepsilon^{(m+n+1)}\mathfrak{R}\left(\hat{\mathbf{V}}^{(m)}\cdot\nabla_{\mathbf{X}}\hat{\Psi}^{(n)^{*}}\right)+\varepsilon^{(m+n)}\mathfrak{I}\left(\hat{\mathbf{V}}^{(m)}\cdot\mathbf{k}\hat{\Psi}^{(n)^{*}}\right)\;.\end{split}$ (20)
Additionally, using the wind amplitude relations derived from the Exner pressure equation (Equation˜11)
$i\mathbf{k}\cdot\hat{\mathbf{V}}^{(0)}$ $=0\;,$ (21)
$i\mathbf{k}\cdot\hat{\mathbf{V}}^{(1)}$ $=-\frac{1}{\overline{P}}\nabla_{\mathbf{X}}\cdot\left(\overline{P}\hat{\mathbf{V}}^{(0)}\right)\;,$ (22)
$i\mathbf{k}\cdot\hat{\mathbf{V}}^{(2)}$ $=-\frac{1}{\overline{P}}\nabla_{\mathbf{X}}\cdot\left(\overline{P}\hat{\mathbf{V}}^{(1)}\right)\;,$ (23)
where $\overline{P}=\overline{\rho}\overline{\theta}$ is the mass-weighted potential temperature of the reference atmosphere. Consequently, Equation˜20 can be rewritten as
$\begin{split}\varepsilon\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{V}}}\right\rangle\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Psi}}\right\rangle=\mathcal{Q}^{(0)}+\mathcal{Q}^{(1)}+O\left(\varepsilon^{4}\right)\;.\end{split}$ (24)
@@
The leading-order terms on the right-hand side of Equation˜20, summarized in Equation˜24 as $\mathcal{Q}^{(0)}$, are $\mathcal{O}(\varepsilon)$ smaller than the large-scale advection and can be rewritten as
$\mathcal{Q}^{(0)}\equiv-\frac{\varepsilon^{2}}{2\overline{P}}\nabla_{\mathbf{X}}\cdot\mathfrak{R}\left(\overline{P}\hat{\mathbf{V}}^{(0)}\hat{\Psi}^{(1)^{*}}\right)\;.$ (25)
@@
Similarly, all terms in Equation˜20 that are $\mathcal{O}\left(\varepsilon^{2}\right)$ smaller than the large-scale advection are grouped into $\mathcal{Q}^{(1)}$ in Equation˜24 and can be rewritten as
$\begin{split}\mathcal{Q}^{(1)}\equiv-\frac{\varepsilon^{3}}{2\overline{P}}&\nabla_{\mathbf{X}}\cdot\mathfrak{R}\left[\overline{P}\left(\hat{\mathbf{V}}^{(0)}\hat{\Psi}^{(2)^{*}}+\hat{\mathbf{V}}^{(1)}\hat{\Psi}^{(1)^{*}}\right)\right]\;.\end{split}$ (26)
The term $\mathcal{Q}^{(0)}$ corresponds to tracer transport by the Stokes drift associated with gravity waves. In contrast, $\mathcal{Q}^{(1)}$ captures higher-order nonlinear contributions, which originate from secondary wave fields generated by the primary (first-order) waves.
@@
With gravity wave parameterization in mind, it is convenient to express the terms in Equations˜25 and 26 using variables commonly computed in parameterization models, such as the wave action density and wavenumbers. This formulation will be presented in the following sections.